OpenStudy (anonymous):

If a bacteria population doubles in 7 days, when was it half of its present population?

OpenStudy (anonymous):

I understand the half of its present population was 7 days ago

OpenStudy (anonymous):

How can you prove it in mathematical form such asA=A_0(1+i)^n

OpenStudy (dumbcow):

\[P_n = P_0(2^{t/7})\] then solve for when P_n = P_0/2 \[\frac{P_0}{2} = P_0 (2^{t/7})\] \[\\2^{-1} = 2^{t/7}\] \[-1 = \frac{t}{7}\] t = -7

OpenStudy (anonymous):

Thank you:)

OpenStudy (dumbcow):

yw